40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 337 of 350  Not logged in ELOG logo
ID Date Author Type Category Subject
  664   Sun Jul 13 22:39:16 2008 JohnSummaryGeneralEdited medm screens
I've edited the FSS and PMC screens so that red boxes are shown around the appropriate slider if a gain or offset is not within the limits defined in C1PSL_SETTINGS_SET.adl

With the current setting of 0 V the FSS input offset is red. According to the settings screen the nominal value is 0.3 +/- 0.050. Are there any objections to editing the nominal value?

I changed the LockMC screen so that red boxes are not shown when the up/down scripts are not running; when they are active you should see a green box.
  663   Sun Jul 13 17:19:29 2008 ranaSummaryPSLMOPA SLOWM Calibration
John, Rana

We first unlocked the FSS and ramped the SLOW actuator. With the PMC locked we observed the PMC PZT voltage
as a function of SLOWM (SLOW loop actuator voltage). We believed this to be ~1-5 GHz / V. Since this is
not so precise we then ran a slow (2 min. period) triangle wave into the slow actuator and looked at the
ref cav transmission peaks to calibrate it.

Plot is attached>

We assume that the reference cavity length = 203.2 mm then the FSR = 737.7 MHz. So looking at the plot
and using our eye to measure the SLOWM calibration is 1054 +/- 30 MHz/V. This is probably dominated by
our eye method.

Note: we tried to get the length from T010159-00-R (Michele, Weinstein, Dugolini). In that doc,
the length used is 203.3 mm whereas its 203.2 mm in the PSL FDD (?). The calculation of the FSR is also
incorrect (looks like they used c = 299460900 instead of 299792458 m/s). We took the length from the PSL FDD
(T990025-00-D) but not the FSR, since they also did not find the right value of 'c'. I guess that the speed
of light just ain't what it used to be.
Attachment 1: SLOWDCcalibration.png
SLOWDCcalibration.png
  662   Sat Jul 12 23:28:31 2008 ranaUpdatePSLPMC needs help
As everyone has noticed recently, the PMC seems to have a PZT problem. It often zooms
of to one of its rails after locking as if the PZT range has decreased dramatically.

WE should check this on Monday by disabling the FSS and applying a slow triangle wave to
the NPRO frequency. The PMC will track this frequency change and this will allow us to
diagnose its problems.

If it has real problems, I have a spare in W. Bridge which we can swap in temporarily.

The attached plot shows 3 years of trend; looks like it went bad in summer of last year.

This also seems to be the cause of our ISS saturation problems: lowering the gain of the PMC
even slightly increased the intensity noise enough to cause saturation. Increasing the gain
even slightly increased the intensity noise enough to cause saturation due to PMC servo oscillation.
Attachment 1: Untitled.png
Untitled.png
  661   Fri Jul 11 23:55:25 2008 alanUpdateLockingRF common mode at zero offset

Quote:
rob, john, yoichi

Last night we succeeded in reducing the CARM offset to zero.



Congratulations! Well done! I look forward to hearing the details and further progress!
  660   Fri Jul 11 20:16:01 2008 EricDAQCamerasTaking data from the GC 750 Camera
Mafalda has been set up with a background process to constantly take data from the GC 750 camera (at the end of the x-arm) for the weekend. This camera will otherwise be inoperable until then.

In the small chance that this slows either Mafalda or the network to a crawl, the process to kill should have PID 26265.
  659   Fri Jul 11 09:29:02 2008 steveUpdatePSLthe PMC still hangs up
Morning alarms:

PMC high voltage is railing, it's auto locker is frosen
I have to move DC Output Adjust slider by hand and it locks immediately

PSL_ISS is saturating. Saturation goes away when PMC is locked.

Laser chiller water is overflowing again: removed 450 cc water
Attachment 1: pmc4d.jpg
pmc4d.jpg
  658   Fri Jul 11 00:30:24 2008 robMetaphysicsComputersstrange SUS controllers

rob, johnnieM

We were hampered early tonight by the fact that someone sneakily turned off the HP RF Ampflier on the AS table.

After that, we were hampered further by mode cleaner strangeness. It would occasionally spontaneously unlock & blow its watchdogs. It never made it through the ontoMCL script (putting DC-CARM onto the MCL). After some investigation, we found that c1susvme1 and c1susvme2 were running stochastically late (SYNC_FE != 0), even though their computation times never got above 61. Also, the end SUS controllers were never late.

Weird.

After rebooting the vertex SUS controllers and the c1lsc, things appear to be working again.
  657   Thu Jul 10 23:27:57 2008 JohnMetaphysicsCamerasSecret handshakes
Rob and I have joined the ranks of the illuminati and exercised our power.


Quote:
Osamu showed me the secret way to change the video labels for the quads and
so we fixed them. He made me swear not to divulge this art.

- Rana Adhikari
  656   Thu Jul 10 19:12:07 2008 AlbertoUpdateGeneralabs cavity length measurement experiment
Yesterday morning, when I started, I found the IFO beam on a different position and the beam spot at the AS port looked very deformed. The overlapping with the secondary beam was not good enough to observe the beats anymore. Restoring the alignments of the interferometers did not work because, as John found out later, some of the photodiodes had offsets and gain which made the restore script ineffective. After resetting the parameters, we had to align every mirror of the interferometers and save the configuration twice. The second times was because on the first time the alignment had been done with the illuminators on. To avoid that in the future, John wrote an alarm to warn about the status of the lights.

After that we fixed the IFO beam, I had to realign the optics in the table to match the secondary beam to the IFO beam. I got the two beam overlapping and, even though the NPRO spot looked distorted, I could observe again some signal of the beat. To do that it was also necessary to have all the interferometer mirrors aligned so that we had more power from the ifo beam although it also made the spot flash. Ideally, to avoid the flashing (which we would also impede the PLL to work) we should work with the interferometer locked. Since that doesn't seem actually possible, we should just keep one of the ITM aligned and improve the beam matching so that we can observe the beats even with less power.

Today I spent the day trying to improve the alignement of the optics to observe the beats with only the ITM aligned, resetting the alignment of both beams with the ireses, with the Farady and all the rest. It was a rather long and tiring process but I think I'm close to the target and maybe tomorrow.
  655   Thu Jul 10 14:59:01 2008 robUpdateLockingRF common mode at zero offset
rob, john, yoichi

Last night we succeeded in reducing the CARM offset to zero.

We handed off control of the common mode servo from PO-DC to POX-I.

We pushed the common mode servo bandwidth to ~19kHz. Without the boosts, it had ~80 degs of phase margin. Didn't measure it after engaging the boosts (Boost + 1 superboost). Trying to engage the second superboost stage broke the lock.

The process is fully scripted, and the script worked all the way through several times.

The DARM ugf was ~200Hz. The RSE peak could clearly be seen. No optical spring, as expected (we're locking in anti-spring mode).

Engaging test mass de-whitening filters did not work (broke the lock).

I'm attaching a lock control sequence diagram and a trend of the arm power during a scripted up-sequence. I think the script can be sped up significantly (especially the long ramp period).

Up next:

Calibrated DARM spectrum
Noise hunting (start with dewhites)
DC - Readout
Lock to the springy side.
Attachment 1: lock_control_sequence_worked.png
lock_control_sequence_worked.png
Attachment 2: trendpowerbuild.png
trendpowerbuild.png
  654   Thu Jul 10 13:47:12 2008 YoichiHowToComputerssvn access via https
Now you can access to the svn repository on nodus by https.
To perform a checkout, you can use the following command

svn co --username svn40m https://nodus.ligo.caltech.edu:30889/svn/trunk/chans

This will check out "chans" directory.
The password for svn40m is written in the usual place.
You can also access the URL by a web browser to see the repository in a very primitive way.
A nice web interface for browsing the repository is planed but not yet implemented.
  653   Wed Jul 9 17:58:19 2008 JohnSummaryGeneralIlluminator alarms
This morning some time was wasted on alignment due to the illuminators.

I added the illuminators to the alarm handler. They will give a RED alarm whenever
they are turned on. You can find the alarms in 40M->Misc->Illuminators.

To do this I edited the Illuminators.db file and restarted c1aux by telneting and typing Ctl-X.
I then added the groups and channels to 40M.alhConfig.
  652   Wed Jul 9 15:04:22 2008 steveMetaphysicsPhotosSURFs helping hands
Surf students are helping out with baffle cleaning.
Attachment 1: surfjob.png
surfjob.png
  651   Wed Jul 9 12:42:14 2008 JohnUpdateLockingHand off to RF CARM
Rob, Yoichi, John

Last night we were able to reduce the CARM offset to around 80. This was achieved by increasing the DARM gain and
switching to AS_I when AS_Q went bad. This is probably a temporary solution, we will probably switch to DC readout
for DARM as we bring the arms on resonance.

Having reduced the arm offset enough to get us into the linear region of the RF_CARM signal (POX_I) we worked on
analogue conditioning of this signal to allow us to hand over. Lock was maintained for over 20 minutes as we did
this work.

We were able to partially switch over both the frequency and length paths to this new signal before losing lock.
Attachment 1: LongLock080709.png
LongLock080709.png
  650   Tue Jul 8 21:58:22 2008 albertoUpdateGeneralSecondaty beam aligned to the IFO beam again
Yesterday the alignment of the secondary beam to the IFO was completely lost and today I had to realign all the optics before I was able to match the two spots again. I had to reset the height of the irises and I had also to replace mirror M1 with one with a larger angular motion. Eventually I obtained the beat again. Working on the optics table I inadvertently misaligned the OSA but I didn't make in time to bring it back before the night shift people came. I'll work on that tomorrow morning.
  649   Tue Jul 8 21:46:38 2008 YoichiConfigurationPSLGC650M moved to the PMC transmission
I moved a GC650M, which was monitoring the light coming out of the PSL, to the transmission port of the PMC to see the transmitted mode shape.
It will stay there unless someone find other use of it.

Just FYI, you can see the picture from the control computers by the following procedure:

ssh -X mafalda
cd /cvs/cds/caltech/target/Prosilica/40mCode
./SampleViewer

Chose 02-2210A-06223 and click on the Live View icon.
  648   Tue Jul 8 12:25:54 2008 JohnSummaryPSLISS gain set to 2dB
  647   Tue Jul 8 10:26:30 2008 MashaUpdateAuxiliary lockingsetup updates
Yesterday I changed one of the beam splitters in the Mach Zehnder to one with a more stable mount as to reduce the system's coupling to environmental noise.

With help from John, I worked out how to get the signals from the two channels of the interferometer into the digital system. I put up BNC cables along part of the Y arm to connect the output of the detectors into the digital channels.
  646   Tue Jul 8 10:20:10 2008 steveUpdatePEMHEPA turned on
It is specially important to run the PSL-HEPA filters when inside counts is peaking at 30,000 counts

There is a small label at the hepa on/off switch:
enclosure open 100%,
low noise off,
normal 60% of Variac voltage setting on the top of the enclosure

Not running the HEPAs will lower the temp fluctuations from 1.5 to o.5 degree C
at the cost of particle counts from 0 to immidiate room counts.
Attachment 1: hepa.jpg
hepa.jpg
  645   Tue Jul 8 08:16:56 2008 steveUpdatePSLPMC problem
The PMC is unhappy. PMC auto locker is not working.
DC output slider adjust has to be moved from rail to rail before it locks.

MZ is working great.
Attachment 1: pmc.jpg
pmc.jpg
  644   Tue Jul 8 00:14:28 2008 JohnSummaryComputersAlarm handler
Rob thought it would be nice to have some alarms on the cpu loads and FE syncs.
I added all these channels to the alarm handler config file and wrote a script
which would set their values (HIHI,HIGH etc).

Ezcawrite allowed me to set the alarm levels (and ezcaread would give the correct
value) but no matter what I set the value to the alarm wouldn't sound.

After experimenting with a few other channels it appears that the alarm handler will
not show alarms if the alarm levels are absent from the db file (even though ezca
gives a value).

I edited the following files so we can have alarms on the cpus.

In c1iscepics:
lsc40m.db
asc40m1.db

In c1losepics:
bs.db
etmx.db
etmy.db
itmx.db
itmy.db
mc1.db
mc2.db
mc3.db
prm.db
srm.db

I saved backups in the appropriate folders.

Next time we have a bootfest please also do c1iscepics and c1losepics so these changes
will be implemented.
  643   Mon Jul 7 19:15:38 2008 AlbertoUpdateGeneralOptics alignement on the ABS length experiment
Today I started setting up the PLL instruments to lock the frequency of the NPRO beam to the IFO beam. with no need of a new alignment after the weekend I was able to see the beat again, although this time I found at a different temperature of the NPRO laser of about 54 degrees (vs 51 of the last time).
I've got the Marconi as local oscillator (LO), the mixer Koji suggetsed, the SR560 and a 5 MHz low pass filter to cut the 33, 66 and 99 MHz present in the output signal from the PD. The filter worked well and I was able to single out only the beat resonance from the power spectrum.
In the attempt to enhance the amplitude of the beat, as Koji suggested, I tried to work on the alignment of the steering mirrors. While I was doing that, for some reason the pre-modecleaner lost the alignment and I had to ask John to help me lock it again. during the process I lost the old alignment but at the end I got a new one, apparently (from the camera) even better than the other. Although after that the beats were gone. Actually after the lock-in of the PMC the IFO beam didn't look as good as before, so it might be also for that reason.

I'll try again tomorrow, after that probably tonight Rob is going to reset the alignments of the interferometers.
  642   Mon Jul 7 16:30:08 2008 steveUpdatePSLPSL-PEM 16 days trend
This morning the laser head temp was up to 20.3C because the laser chiller was overflowing.
I removed 700 cc water.

The PSL-FSS_RMTEMP became much more stable during the holidays as the psl enclosure was closed for 4 days

The high particle counts can be explaned by construction activity today.

The PMC & MZ PZT high voltages were out of range this morning.
Attachment 1: pempsl16d.jpg
pempsl16d.jpg
  641   Mon Jul 7 14:02:05 2008 YoichiUpdateComputersSVN conversion progress
So far /cvs/cds/caltech/medm, /cvs/cds/caltech/chans and /cvs/cds/caltech/scripts have been converted to svn working copies.
Now /cvs/cds/caltech/target is being converted.
  640   Mon Jul 7 13:58:37 2008 Eric, josephbDAQPEMUsing unused PEM channels to test camera code
Joe and I have taken control of the EPICS channels C1:PEM-Stacis_EEEX_geo and C1:PEM-Stacis_EEEY_geo since we heard that they are no longer in use.  We are currently 
using them to test the ability for the Snap camera code to read and write from EPICS channels.  Thus, the information being written to these channels is completely unrelated
to their names or previous use.  This is only temporary; we'll create our own channels for the camera code shortly (probably within the next couple of days).

- Eric
  639   Mon Jul 7 13:49:27 2008 YoichiHowToPSLMZ offset, gain tips
John, Yoichi

This morning John found that MZ servo is not working.
We were able to bring the MZ back by changing the output offset a bit. But we were not sure what was actually wrong.
So we pulled out the MZ board and checked several TPs to understand the behavior.
Here is the summary of what we learned this morning.

The MZ control board has the following stages:

[Mixer] -(error signal)-> [Sum amp for input offset] -(error + offset)-> [Variable Gain Amp] -> [Filter (x100 DC gain)] -(FB signal)-> [High Voltage Amp] -> output
(The HV amp also works as the sum amp for the output offset)

(1) We noticed that the Sum amp for the input offset has an output of -0.14V even when the offset input is 0V. This can be canceled by the input offset slider.
So for the moment, it is fine. But we might want to change the op-amp because the weird offset implies there might be something wrong with the chip.
The procedure to null the -0.14V offset is the following:
a) Enable Test 1 input on the MZ MEDM screen.
b) Move the input offset slider until the mixer monitor becomes 0V. Currently the input offset slider is at -7.5V to cancel the -0.14V offset.

(2) Because the gain of the Variable Gain Amp and the Filter combined is large, the Filter can be easily saturated if the output offset is not right.
This was the cause of the MZ problem this morning. The output offset slider was at a wrong position making the error signal slightly off centered from zero.
This residual DC error signal was amplified by the large gain chain and saturated the filter amp.
Our experience is that the output offset cannot go below -3V. We set it at 0V for now.
  638   Mon Jul 7 13:06:38 2008 KojiUpdateGeneralBeats of the two lasers in the absolute length measurement observed
One may need an RF filter after the mixer. I expect the SR560 does work for this purpose.
If it does not, a passive LPF can be used.


Quote:
I didn't post a screenshot from the RF SA because I had troubles with the interface with the computer (unfortunately the network SA cannot export the data).

There's is problem with the PLL circuit. The signal, beside the beat, also contains peaks at 33, 66 and 99 MHz, so we should think about filtering those out, correct?
  637   Mon Jul 7 11:22:02 2008 AlbertoUpdateGeneralBeats of the two lasers in the absolute length measurement observed
I didn't post a screenshot from the RF SA because I had troubles with the interface with the computer (unfortunately the network SA cannot export the data either).

There is problem with the PLL circuit. The signal, beside the beat, also contains peaks at 33, 66 and 99 MHz, so we should think about filtering those out.


Quote:
Great! Conguraturation! I wish if I could see it! It's nice if you can put the photo or anything of the RF spectrum analyzer.

Next step:
o You can try to maximize the beat amplitude by the tuning of the Injection steering mirrors.

o At the south end of the SP table, I prepared a frequency mixer. You can put the beat signal into the RF input, and an oscillator (which you can bring from somewhere) to the LO input in order to obtain the error signal of the PLL. Put the IF output of the mixer in a SR560, and please try to lock it by a simple 6db/oct (1st order) LPF of the SR560. For the actuator you can use the fast-pzt input of the NPRO.


Quote:
Then I explored the range of temperature of the NPRO from 35deg (C) to 51.2807deg and at that point I could observe a peak corresponding to the beat at about 10MHz on the network analyzer.
  636   Sun Jul 6 16:17:40 2008 tobinHowToComputersSVN
I was able to check out the 40m SVN here in Livingston using this command:

svn co svn+ssh://controls@nodus.ligo.caltech.edu/cvs/cds/caltech/svn/trunk/medm

As you might guess, this uses ssh in place of the web server (which we don't have yet).
  635   Thu Jul 3 22:54:45 2008 KojiUpdateGeneralBeats of the two lasers in the absolute length measurement observed
Great! Conguraturation! I wish if I could see it! It's nice if you can put the photo or anything of the RF spectrum analyzer.

Next step:
o You can try to maximize the beat amplitude by the tuning of the Injection steering mirrors.

o At the south end of the SP table, I prepared a frequency mixer. You can put the beat signal into the RF input, and an oscillator (which you can bring from somewhere) to the LO input in order to obtain the error signal of the PLL. Put the IF output of the mixer in a SR560, and please try to lock it by a simple 6db/oct (1st order) LPF of the SR560. For the actuator you can use the fast-pzt input of the NPRO.


Quote:
Then I explored the range of temperature of the NPRO from 35deg (C) to 51.2807deg and at that point I could observe a peak corresponding to the beat at about 10MHz on the network analyzer.
  634   Thu Jul 3 18:48:09 2008 AlbertoUpdateGeneralBeats of the two lasers in the absolute length measurement observed
I adjusted the alignment of the flipper mirror as suggested by Koji making the two beam spots match. I also aligned all the IFO mirrors (ITMs, PRM, SRM, ETMs) to have more power for the IFO signal at the AS port. When I did that I could see the beats at the AS OSA. Then I explored the range of temperature of the NPRO from 35deg (C) to 51.2807deg and at that point I could observe a peak corresponding to the beat at about 10MHz on the network analyzer. The peak tends to drift because the laser takes probably a longer time to actually thermalize and it moves very rapidly changing the temperature of the laser.
  633   Thu Jul 3 16:57:23 2008 JohnSummaryGeneralFSS_RMTEMP
The FSS room temp alarm has been beeping a lot recently. I altered the FSS_RMTEMP alarm levels using the
same method as Rana.

The alarm is still souding so, at least by my calculations, it must be colder than usual.
Attachment 1: FSStime.png
FSStime.png
Attachment 2: FSSalarm.png
FSSalarm.png
  632   Thu Jul 3 16:18:51 2008 robSummaryLockingspecgrams
I used ligoDV to make some spectrograms of DARM_ERR (1), QPDX (2), and QPDY (3). These show the massive instability from 30-40Hz growing in the XARM in the last two minutes of a reasonably high power lock (arm powers up to 30). It's strange that it only shows up in one arm.

CARM is on PO-DC, for both the MCL and the AO path.
DARM is on AS166Q.
Attachment 1: darm_specg.png
darm_specg.png
Attachment 2: qpdx_specg.png
qpdx_specg.png
Attachment 3: qpdy_specg.png
qpdy_specg.png
  631   Thu Jul 3 13:54:26 2008 robConfigurationComputersmDV on rosalba

Does mDV work on rosalba? It can't find NDS_GetChannels. Looking on mafalda, I see that NDS_GetChannels is a mexglx. I think this means someone may need to compile it for 64-bit matlab before we can have mDV on rosalba. When that's done, we should get mDV running on megatron.
  630   Thu Jul 3 13:12:32 2008 Rob, Yoichi, JohnUpdateLockingMore oscillations
Bounce/ roll filters were added to the short degrees of freedom to reduce the effect of the 24Hz line seen on Tuesday night.

However last night saw the arrival of a new oscillation at ~34Hz. This may be the second harmonic of the MOS roll mode. Reducing the arm offset would cause this oscillation to ring up and break the lock (first plot). This effect was repeatable.

No signal was seen in the oplevs or osems which leads us to rule out alignment problems, at least for now.

Although one can clearly see DARM_ERR increasing as arm power increases adding a resonant gain in the DARM loop had no effect.

We also noticed that x arm transmission was significantly more noisy than the Y (second plot). And showed greater coherence with the increase in DARM noise. Investigations showed that the PD was not the source of the difference.

Turning up the MC gain seems to help a little.

We're now looking at POX as a candidate for RF_CARM (third plot).
Attachment 1: LOL.png
LOL.png
Attachment 2: NoisyX080702.png
NoisyX080702.png
Attachment 3: POXforCARM080702.png
POXforCARM080702.png
  629   Thu Jul 3 12:36:05 2008 JonhSummarySUSETMY watchdog
ETMY watchdog was tripped. I turned it off and re-enabled the outputs.
  628   Thu Jul 3 11:53:30 2008 KojiUpdateGeneralStatus of the alignment of the NPRO beam for the Absolute Length Measurement
The method itself looked fine.

Use of the one arm configuration will make the work easier as constant power at the AS port is obtained.
How much is the FSR of the OSA?

Apparently the alignment is not good any more as Alberto pointed. Everytime you touch the flipper you'd better to adjust it. Then, if necessary, adjust the injection steerings.

If the PSL beam is blocked, only the injection beam appears at the optical ports. The spot is obtained at the AS port and the SY port (REFL) at the same time. I recommend to confirm the transmittion to the SY port too by the CDD, the card, and whatever. Note that this may be difficult because this will have the beam power of below 1 mW.


Quote:
Alberto> Today I've tried to bring the frequency of the NPRO laser ...
  627   Wed Jul 2 19:15:52 2008 AlbertoUpdateGeneralStatus of the alignment of the NPRO beam for the Absolute Length Measurement
Today I've tried to bring the frequency of the NPRO laser close enough to that of the IFO beam so that the beat between the two beams can be at a detectable frequency for the photodiode. The way I've been changing the frequency is by the NPRO's temperature control on its driver.

Looking at the signal from the AS OSA should enable us to monitor the direction in which the frequency is changing. Every time the resonances of the IFO beam and of the NPRO beam overlap, we know that the frequencies of the two beams are some FSR of the OSA away from each other. At the overlapping of the resonances, if the difference of frequency is within the detectable range of the photodiode, we should see a peak in the network/spectrum analyzer.

This way turned out not very easy in practice because from the AS OSA one can hardly distinguish the resonances of the primary beam from those of the secondary beam. The cause is mainly the flashing of the IFO beam at the AS port which produces a pattern of resonances of different amplitude. Also for some reason, triggering the output signal from the OSA at the oscilloscope doesn't work very well.

However, even if we didn't have these problems, I think that the two beams are not very well aligned, at least not anymore. I'm attaching some pictures from the AS port. The bright spot on the left is the NPRO beam and the one in the center which flashes is the IFO beam. We probably need some more work in the alignment of the NRPO beam.
Attachment 1: DSC_0156.JPG
DSC_0156.JPG
Attachment 2: DSC_0158.JPG
DSC_0158.JPG
  626   Wed Jul 2 18:30:01 2008 JohnUpdateIOOQPD alignment
I aligned the beams on the following QPDs

IP POS
IP ANG
MC TRANS
IOO POS
Attachment 1: IOO080702.png
IOO080702.png
Attachment 2: MC080702.png
MC080702.png
  625   Wed Jul 2 17:19:03 2008 JohnSummaryComputersop440m - shutdown and restarted
After 160days op440m was getting a little slow.
  624   Wed Jul 2 15:14:42 2008 steveUpdateGeneraladded beam traps
I placed baked razor beam trap after INJ_SM1 and flipper in the injection path on the AP table


Quote:
I have constructed the beam injection optics for the abs length measurement.

The injection beam was coarsely aligned to the interferometer. The reflected beam from SRM was already seen at AS CCD.
I have attached the optical configration for this measurement and the optics layout at the AP table.

I am going to go to LHO for three weeks. During the absence Alberto tunes the mode matching and the alignment of the interferometer.

In the process of making this report, I noticed that one of the iris apertures is about disturbing the beam for OMCR CCD. I will check this before I go to Hanford. Also an RF spectrum analyzer is at the AP table. I try to return this near the PSL on Monday morning.

Attachment 1: Optical configuration for the abs length measurement.
1) One of the arms is locked to the PSL beam by the main control system (red).
2) A laser beam is injected from the AS port (blue). This laser essentially has different frequency from that of PSL.
3) The injected beam and the outgoing PSL beam appear at the output of the faraday in the injection system.
4) They beat each other at the frequency difference of those two lasers.
5) A PLL is used to lock the frequency difference to a local oscillator (LO).
6) The LO frequency is swept at around 3.87MHz, that is the approximate FSR frequency of the arm cavity.
7) If the LO frequency hits the FSR within the resonant width, the beating also appears at the transmitted light as the injected beam also becomes resonant to the arm cavity.
8) Amplitude of the beating at the transmitted light is measured by a RF spectrum analyzer as a function of the LO frequency. We get the FSR frequency (= the arm cavity length) from the top of the resonance.

Attachment 2: Optics at the AP table for the laser injection
700mW NPRO, laser source. vertically polarized.
Periscope, to raise the beam 1 inch to make the beam at the 4 inch elevation.
INJ_SM1/INJ_SM2, steering mirrors to align the injection beam to the IFO beam.
HWP1, half wave plate to make the beam to the farady horiz-polarized. nominal 42deg on the readout.
FI, Faraday isolator for protection of the NPRO from the returning light, for obtaining the returning light.
HWP2, to make the beam from the Faraday horiz-polarized. nominal 357deg on the readout.
MM_Lens, f=125mm to match the laser mode to the IFO beam.
SM1/SM2, steering mirrors to align the IFO beam to the Farady Isolator.
IRIS1/IRIS2, for the coarse alignment of the injection beam.
FLIP, flipper mount to turn on/off the injection optics.

Alignment procedure of the injection system
0) Ignite NPRO several hours before the experiment so that the laser frequency can be stable.
1) Turn up FLIP. Close the shutter of NPRO.
2) Adjust SM1/SM2 so that the ifo beam can appear at the output of FI.
3) Adjust height and position of IRIS1/IRIS2 with regard to the ifo beam so that the ifo beam goes through IRIS1/IRIS2 even when they are closed.
4) Turn down FLIP. Open the shutter of NPRO.
5) Adjust INJ_SM1/INJ_SM2 so that the injection beam can go through IRIS1/IRIS2 even when they are closed.
6) At this time, it is expected that the reflection of the injection beam from SRM appears at AS CCD, if SRM is aligned.
7) Adjust INJ_SM1/INJ_SM2 so that the injection beam at AS CCD can overlap to the IFO beam.
8) Confirm the beam at the output of the FI also overlaps.
---- We are here ----
9) Change the ifo configuration to the X or Y arm only.
10) Scan the crystal temperature of the 700mW NPRO in order to try to have the beating of the two beams at the PD. AS OSA may be useful to obtain the beating.
11) Once the beating is obtained, adjust INJ_SM1/INJ_SM2 such that the beating amplitude is maximized.
  623   Wed Jul 2 13:56:10 2008 Rob, Yoichi, JohnUpdateLocking24.5 Hz resonance
Work continues on trying to reduce the CARM offset using dc signals from PO_DC. Got up to arm powers of
~35 last night.

We found that progress was stymied by an oscillation around 24 Hz. This oscillation was clearly visible
in the intensity of the light at REFL, PO and TrX.

Initially we suspected that this oscillation was due to an instability in the CARM loop. We attempted to
solve the problem by tuning the crossover frequncy of the AO and MC_L paths and shaping the MC_L loop to
reduce the impact of the 24 Hz noise.

After some quick tests we found that the 24 Hz signal was present even when dc CARM was used. It appears
that the peak is in fact due to a SOS mechanical resonance. We currently suspect a roll mode.

We're going to check that PRC, MICH and DARM have filters to attenuate the 24 Hz line. We'll also look at the
SUS_POS bandstop filters to see where they are centred.

The ISS was behaving strangely again. Constantly saturated at 5dB of gain. Someone needs to look a this.
Attachment 1: locking080702.png
locking080702.png
  622   Wed Jul 2 10:35:02 2008 EricSummaryCamerasGeneral Summary
I finished up the 2D Gaussian fitting code, and, along with Joe, integrated into the Snap software so that it automatically does a fit to every 100th image. While the fitting works, it is too slow for use in any feedback to the servos. I put together a center of mass calculation to use instead that is somewhat less accurate but much faster (almost instantaneous versus 5-10 seconds). This has yet to be added to the Snap software, but doing so would not be difficult.

I put together a different fitting function for fitting the multiple lorentzian resonance peaks in a power spectrum that would result from sweeping the length of any of the mode cleaners. This simply doesn't work. I tested it on some of Josh Weiner's data collected on the OMC last year, and the data fits poorly. Attempting to fit it all at once requires fitting 80000 data points with 37 free parameters (12 peaks at 3 parameters per peak and 1 offset parameter), which cannot be done in any reasonable time period. Attempting to fit to one specific peak doesn't work due to the corruption of the other nearby peaks, even though they are comparatively small. The fit places the offset incorrectly if given the opportunity (green line in attemptedSinglePeakFitWithoutOffset.tiff and attemptedSinglePeakFitWithoutOffsetZoomed.tiff). Removing this as a parameter causes the fit to do a much better job (red line in these two graphs). The fit still places the peak 0.01 to the right of the actual peak, which worse than could simply be obtained by looking at the maximum point value. Additionally, this slight shift means that attempting to subtract out the peak so that the other peaks are accessible doesn't work -- the peaks are so steep that the error of 0.01 is enough to cause significant problems (red in attemptedPeakSubtraction.tiff is the attempted subtraction). Part of the problem is that the peaks are far from perfect lorentzians, as seen by cropping to any particular peak (OMCSweepSinglePeak.tiff ). This might be corrected in part by correcting for the conversion from PZT voltage to position, which isn't perfectly linear; though I doubt it would remove all the irregularities. At the moment, the best approach seems to be simply using a center of mass calculation cropped to the particular peak, though I have yet to try this.

Changing Josh's code to work for the digital cameras and the PMC or MC shouldn't be difficult. Changing to the MC or PMC should simply involve changing the EPICs tags for the OMC photodiodes and PZTs to those of the PMC or MC. Making the code work for the digital cameras should be as simple as redirecting the call to the framegrabber software to the Snap software.
Attachment 1: attemptedSinglePeakFitWithoutOffset.tiff
Attachment 2: attemptedSinglePeakFitWithoutOffsetZoomed.tiff
Attachment 3: attemptedPeakSubtraction.tiff
Attachment 4: OMCSweepSinglePeak.tiff
  621   Wed Jul 2 06:46:05 2008 AlbertoConfigurationGeneralNPRO on to warm up
This morning I turned on the NPRO on the AP table so that it can warm up for a few hours before I start using it today.
The flipping mirror is down so no beam is injected in to the IFO.


Alberto
  620   Wed Jul 2 00:59:13 2008 MashaUpdateAuxiliary lockingbalanced detection, noise plots, progress
Progress report submitted today(!). It is on the 40m wiki page. Below is a figure of some estimated noise sources.

Made voltage divider that acts as an attenuator for one of the paths in the Mach Zehnder, which should help to balance the detection and reduce noise.

First tested using a 636 Hz Matlab generated audio signal (thanks to John inspiration on portable headphones). Figure is attached, with plots of noise spectra of original and optimized signal with and without added acoustic noise (visible as peaks as 636 and 1272 Hz, linewidth approx 4 Hz). My first try at optimization reduces the noise by nearly an order of magnitude for most frequencies.

Will work on finding different noise source to better see what happens at low frequency, and try to get finer control of tunable gain.
Attachment 1: noise_sources.pdf
noise_sources.pdf
Attachment 2: balancing_detectors.png
balancing_detectors.png
  619   Tue Jul 1 21:54:05 2008 KojiUpdateGeneralRe: Abs. Length Meas. setup
I tried to look for the beating in the signal from the PD but I couldn't find. I had the temperature of the laser initially set to 40deg and then slowly increased by one degree. The manual of the laser says the frequency should change by several GHz. The problem is then that our PD is limited to no more than 30Mhz.

Although the two beams seem to overlap quite well, we might still need a better matching of the injected beam.


Alberto


Quote:
o The position of the iris was adjusted so as not to disturub the beam for OMCR CCD.

o The RF spectrum analyzer was returned to the place of the network analyzer.


Quote:

In the process of making this report, I noticed that one of the iris apertures is about disturbing the beam for OMCR CCD. I will check this before I go to Hanford. Also an RF spectrum analyzer is at the AP table. I try to return this near the PSL on Monday morning.
  618   Tue Jul 1 21:45:48 2008 JohnUpdatePSLMach Zehnder script and screen
I've edited C1: PSL_MACH_ZEHNDER.adl and /cvs/cds/caltech/scripts/PSL/MZ/lockMZ
to reflect the changes described in entry #616.
  617   Tue Jul 1 21:27:27 2008 robHowToComputer Scripts / Programsslider twiddling after reboot

Sometimes after we reboot the front-end machines, some of the hardware gets stuck in an unknown state. We generally fix this by twiddling EPICS settings, which refresh the hardware somehow and put it into a known state. I've started a script (slider_twiddle) which we can just run after reboots to do this for us. Right now it just has the QPD whitening gain settings. As we find more stuff, we can add to it. It's in $SCRIPTS/Admin/.
  616   Tue Jul 1 16:48:42 2008 rob, johnConfigurationPSLMZ servo switch problem resolved forever

Quote:
C1:PSL-MZ_BLANK switch (to turn on/off the servo) is not working again. The switch is always off regardless of the epics state.
I pushed the cables into the xycom card, but it did not fix the problem.


We have fixed this problem forever, by totally disabling this switch. Looking at the schematic for the MZ servo and the datasheet of the AD602, we found that a HI TTL on pin 4 disables the output of the AD602. Since the MZ servo was stuck in the off position, this seemed to indicate that it may be the XYCOM220 itself which is broken, constantly putting out a +5V signal regardless of the EPICS controls. We thought we might be able to get around this by disconnecting this signal at the cross-connect, but ultimately we couldn't find it because there is no wiring diagram for the Mach-Zehnder (!). So, we pulled the board and wired pin 9A of P1 to ground, permanently NORMALizing the MZ_BLANK switch. John has marked up the schematic, and someone should modify the MEDM screen and check the new screen into svn.

We can still the turn the MZ servo on and off by using the test input 1 switch.

Someone also will need to modify the MZ autolocker to use the test input 1 (MZ_SW1) instead of the old MZ_BLANK.
  615   Tue Jul 1 14:24:58 2008 robHowToComputer Scripts / Programsconlog time machine

I've written a perl script (now in the $SCRIPTS/general directory) which implements a "conlog restore" command, restoring channels matching a regexp to a given time using the conlog records and the EpicsTools.pm perl module. The script is called time_machine_conlog:


Quote:


op440m:~>time_machine_conlog

time_machine_conlog restores EPICS control settings using a conlog time
usage: time_machine_conlog [<--dryrun>] <date=yyyy/mm/dd,hh:mm:ss> <timezone> <regexp>

Can also accept a gps time, in which case timezone=gps.
Use the option <--dryrun> to see conlog output without restoring any settings.

EXAMPLE: time_machine_conlog 2008/05/30,12:00:00 PDT "C1:SUS-MC.*_(PIT|YAW)_COMM"



It sometimes returns an error message even when the command is successful--this is because conlog stores EPICS settings to an absurd level of precision, but ezcawrite will not write EPICS values to this level (or at least won't indicate if it did). I consider this a bug in ezcawrite so I'm not touching it.

The script is untested with regards to switch settings (such as ENABLE/DISABLE). It's mainly intended for numerical values.
ELOG V3.1.3-